# Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval

### v1

@@ 3.3 Zero-Shot Text-to-Video Retrieval
MSR-VTT (Xu et al., 2016) is one of the well-known benchmarks. However, the queries in the dataset can tend to be more generic, e.g., “a tornado and a thunderstorm are looming in the background,” rather than targeting rich information about a specific event, like “2025 LA Fire” from Fig. 1. In contrast, MultiVENT (Sanders et al., 2023) defines a multilingual video retrieval task where the queries refer to specific events.
4 Q2E: Query-to-Event Decomposition
Figure 2: The complete method of our text-to-video retrieval system. The top blue box represents the query decomposition module of our method, while the bottom green box illustrates the video captioning and audio transcription module. The left orange box represents our multi-layer ASR module. The right purple box fuses the ranks calculated from query-video, query-descriptions, and event-descriptions (event = prequel + current + sequel, description = multimodal description). Text-to-text similarity scores were computed using ColBERT (Khattab and Zaharia, 2020), and text-to-video similarity scores were computed using MultiCLIP (Delitzas et al., 2023). In the non-ASR variant of our method, the components within the orange box and orange dotted lines are excluded.
Modern dense retrieval approaches commonly compute an embedding-based similarity (e.g., cosine) between a query and each candidate; those scores rank candidates, and the top-scoring candidates are returned. For example, in video retrieval, one could embed the user’s query using an appropriate LLM, embed the video using an appropriate video embedder, and then compute the cosine similarity between those embeddings Sanders et al. (2023). This “query vs. video” similarity score is used as a proxy for how relevant to video is to the query.
Our approach (shown in Fig. 2) builds upon this general embedding similarity computation, though to better capture the diverse and nuanced multimodal information conveyed in videos and human queries—including the event decompositions Q2E produces—we must extend those similarity computations. In order to capture the rich event information we get from our decomposition (§ 4.1), a single embedding representing the video is not sufficient. We need to additionally get multimodal descriptions of the video (§ 4.2 and 4.3); these are text-based descriptions of the video and audio in the video, which can be more readily evaluated against the query. Using these components, Q2E computes five types of scores $S_{i,v}$ for each video $v$: (a) a query vs. video score, as described above; (b) an event decomposition vs. multimodal description score for each of the prequel, current, and sequel decompositions; and (c) a query vs. multimodal description score. While (a) helps to identify coarse-grained features such as “fire” or “flood”, both (b) and (c) provide fine-grained contextual details. By aggregating these components using entropy-based fusion (§ 4.4), they complement each other to improve the retrieval of specific video events.
4.1 Decomposing Query Events
Like any traditional text-video retrieval system, a user provides a natural language query, which can be a couple of words to 10+ words. In this method, we have used an LLM to decompose it into three sub-events: (1) Prequel, (2) Current, and (3) Sequel. Based on initial experiments, and to manage computation, we only considered five each of prequel, current, and sequel events. We provide our template prompts for each in § A.4.
Prequel
“Prequel” primarily refers to events that can lead to the current event described in the query. An event can have multiple prequels; for example, a wildfire can have strong wind, high-temperature alert, and dry lightning as prequels. Prequel events may enable the event in the query, but do not necessarily cause it.
Current
Sequel
“Sequel” events can result from the event described in the query. Like the prequel, one event can lead to multiple sequel events.
Decomposition Refinement
We refine these event decompositions with temporal, spatial, and primary event data. For this, we have used the same LLM to extract the temporal (2025), spatial (LA), and primary event (Fire) information. Although this rich information can be combined templatically, e.g., “P/S/C (2025, LA, Fire),” during initial development experiments, we found that refining it to read as a more natural, human-provided query markedly improved performance. We produced this refined query automatically using an LLM; we show this in § A.4.3, and provide all responses as part of our anonymized data release. [^2] https://github.com/dipta007/Q2E
4.2 Video Decomposition
In practice, we found that embeddings of videos / video frames were not sufficient to capture the varied actions described by our event decomposition through basic similarity computations. While more advanced similarity approaches, such as cross-encoders, may be feasible, they present computational challenges (requiring running the cross-encoder on every video/query pair). To address this, we instead considered how to extract text descriptions (captions) from the videos themselves, with the aim that these descriptions would be sufficiently similar to our query event decomposition. We found two approaches to be effective for capturing complementary local and global elements across the video: (1) generating contextualized descriptions per selected frame, and (2) generating holistic video descriptions.
Contextualized Frame Descriptions
We sampled K=16 frames from the videos (determined by trading off performance vs. computation—see ablation in § A.1.2). We have tried uniform sampling and a scene-change-based sampling method  [^3] https://www.scenedetect.com/ for frame sampling, and found that uniform sampling, despite its simplicity, works better than scene-change-based sampling (ablation in § A.1.3). As the video contains temporal information and by design, video frames should be connected semantically, so rather than captioning each frame in isolation, we used a sliding window (window size=1) approach and provided the VLM model with the previous frame’s caption (when available). Conditioning on previous frame’s captions helped contextualize subsequent captions. Per video, we have K contextualized frame captions.
Video Descriptions
We found that current VLM models can perform dense captioning but miss the “bigger picture.” For instance, a VLM might describe a video of a wildfire with people running and fire everywhere but fail to recognize the higher-level conclusion of “wildfire.” To address this limitation, we provided an LLM model with all 16-frame descriptions from above and asked it to summarize them into a single video-dense caption, preserving the order of the frame captions.
4.3 ASR Augmented Video Captions
Speech is a critical component of many videos, especially when visual elements alone may not be sufficient enough to convey the full context. Despite using a multilingual ASR model (Radford et al., 2023), the multilingual videos posed a unique challenge to get consistent and accurate ASR transcriptions. To address this, we propose a multi-layer translation pipeline (Fig. 2, left orange box). This pipeline leverages three models: (1) a multilingual ASR model, whisper-v3 (Radford et al., 2023), which transcribes audio into both the original language and its English translation; 2) a translation model, NLLB (Costa-jussà et al., 2022) to generate English transcriptions by translating the original transcript; and (3) an LLM, Llama-70B (Dubey et al., 2024), which refines the English translations produced by the previous two models. Early experiments showed all three were critical for sufficient transcription and translation quality.
4.4 Scoring and Score Fusion
Dense Video Embeddings
We utilized the original query and video to compute a similarity score. The MultiCLIP (Delitzas et al., 2023) model was selected due to its superior performance with multilingual videos compared to the CLIP (Radford et al., 2021) model. Although MultiCLIP is traditionally designed for images, we adapted it for video by following the approach of Sanders et al. (2023): uniformly sampling 16 frames and averaging their embeddings to generate the final video embedding. Cosine similarity was applied to compute the similarity score, which was then scaled to the 0–100 range to ensure comparability with other scores: $S_{i,v}|Q=100.0\times\operatorname{multiCLIP}(Q,V),$ where $S_{i,v}|Q$ denotes the score vector for a given query $Q$, and $\operatorname{multiCLIP}(x,y)$ returns the similarity score between x and y.
[…]
MultiVENT
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
MultiCLIP + ASR 10.08 46.64 74.70 90.73 84.32 68.57 0.93 78.76 88.87 22.66 6
MultiCLIP + Q2E 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
MultiCLIP + Q2E + ASR 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
[…]
InternVideo2-1B + ASR 9.68 40.48 62.43 87.64 73.59 57.34 0.91 68.30 83.84 109.70 7
InternVideo2-1B + Q2E 9.54 40.88 63.40 87.26 75.21 58.73 0.92 69.15 83.96 53.84 7
InternVideo2-1B + Q2E + ASR 10.24 44.94 70.79 92.28 81.85 65.14 0.95 76.10 88.09 42.81 6
MSR-VTT-1kA
MultiCLIP 43.52 69.05 76.88 43.62 13.85 7.71 0.54 59.72 54.27 20.29 2
MultiCLIP + ASR 45.43 70.75 77.19 45.53 14.19 7.74 0.56 61.22 56.11 31.31 2
MultiCLIP + Q2E 44.52 71.26 79.40 44.62 14.29 7.96 0.56 61.51 55.84 17.91 2
MultiCLIP + Q2E + ASR 46.23 73.37 81.71 46.33 14.71 8.19 0.58 63.59 57.83 18.73 2
[…]
InternVideo2-1B + ASR 53.17 73.27 81.61 53.27 14.69 8.18 0.62 67.08 62.49 23.41 1
InternVideo2-1B + Q2E 53.47 73.57 82.11 53.57 14.75 8.23 0.62 67.16 62.47 16.73 1
InternVideo2-1B + Q2E + ASR 56.28 76.58 83.72 56.38 15.36 8.39 0.65 69.53 65.06 15.85 1
Table 1: Performance comparison on MultiVENT and the MSRVTT “1k-A” split. We use MultiCLIP (per Sanders et al. (2023)) and InternVideo2 (the best-performing method on the leaderboard) as our baseline encoders. Our method (Q2E), even without ASR, improves NDCG by 5–19 points on the MultiVENT dataset and 1-2 points on the MSRVTT dataset. Adding ASR further enhances performance, with an additional 3–7 points on MultiVENT and 2 points on MSRVTT, resulting in a total improvement of approximately 8–26 points on MultiVENT and 3–4 points on MSRVTT, across two different encoders.
Query vs. Multimodal Descriptions
In this approach, we used the original query but replaced the video content with the Multimodal Description Set, yielding a text-to-text similarity score instead of text-to-video. Specifically, we compute the maximum similarity between the query and all multimodal descriptions. Formally:
$S_{i,v}|Q=\max_{c=1:C}Sim(Q,Cap(v)_{c}),$ (1)
where $C$ is the number of descriptions associated with the video, and $Cap(v)$ returns the multimodal description of the video $v$. $Sim(x,y)$ computes the ColBERT444 We experimented with cosine-based SBERT (Reimers and Gurevych, 2019) and found that ColBERT outperforms SBERT in both accuracy and speed. Notably, in our case, the captions generated by LLMs may have some noise. ColBERT mitigates the impact of this noise by leveraging maximum aggregation at the token level, rather than compressing information into a single vector. This approach effectively reduces or ignores noisy elements in the multimodal descriptions. Additionally, prior studies (Warner et al., 2024) have shown that ColBERT’s approach minimizes information loss introduced by the averaging process used in SBERT, further enhancing its performance. similarity score between $x$ and $y$.
Event Decomposition vs. Multimodal Descriptions
We use the refined prequel/current/sequel events from the query and the multimodal description set to compute a many-to-many similarity score. Specifically, we compute the maximum similarity score across all descriptions for each event and subsequently take the maximum similarity score among all the events as the final score. This approach reduces the negative impact of a single noisy decomposition or caption by leveraging the global maximum instead of averaging, which is more vulnerable to noisy data. Formally:
$S_{i,v}|Q=\max_{l=1:L}\max_{c=1:C}Sim(E_{i}(Q)_{l},Cap(v)_{c}),$ (2)
4.4.1 Score Fusion
As Q2E is zero-shot, fusing all the scores presents a challenge. Naively averaging all scores would assume equal importance for each score, which does not reflect real-world scenarios. Different scores contribute differently to the final result depending on the query, e.g., for videos relating to earthquakes, the Sequel vs. Descriptions score carries significantly more weight than the Prequel vs. Descriptions score, as there is no visual indication of an earthquake before it occurs. Yin and Jiang (2024) demonstrated that entropy-based methods outperform others in zero-shot rank fusion tasks. In this entropy ranking, disparate similarity scores for an individual video are normalized to a softmax distribution, where low entropy indicates high confidence, while high inverse entropy corresponds to high confidence. Building on this insight, we use inverse entropy rank fusion to combine the scores.
$\hat{S}=\sum_{i}^{5}\frac{1}{H(P_{i})}*P_{i}.$ (3)
By applying Eq. 3, we effectively assign greater weights to scores with lower entropy, emphasizing more confident predictions. This $\hat{S}$ contains an aggregate score for each video; $\hat{S}$ can be sorted to return the ranked list of videos for the query.
5 Results
5.1 Datasets and Implementation
We experimented with the event-centric MultiVENT dataset (Sanders et al., 2023) and the widely-used MSR-VTT dataset (Xu et al., 2016). MultiVENT provides enriched, event-specific human queries, while MSR-VTT contains 10,000 video clips across 20 categories, with each clip annotated with 20 English sentences. We used the standard “1k-A” test split of 1,000 videos for computational constraint, as used in prior works (Jiang et al., 2022; Guan et al., 2023; Jiang et al., 2023a). Dataset details are in Table 13. Full implementation details can be found in § A.3, though briefly, we used Llama-3 as the LLM, InternVL-2.5 as the VLM, and Whisper-large-v3 as the multilingual ASR models. Additionally, we used ColBERT and MultiCLIP for similarity score calculations.
5.2 Zero-Shot Text-to-Video Retrieval
The comparison of our method with the baseline for MultiVENT and MSR-VTT-1kA is shown on the Table 1. We use the standard retrieval metrics: recall/precision at rank K, mean reciprocal rank (MRR), mean average precision (MAP), normalized discounted cumulative gain (NDCG), and mean and median rank (MnR, MdR).
Overall, combined with ASR, Q2E gives the best score in all the metrics across two different datasets and two different text-to-video encoders. But even without audio information, we outperformed the baseline in all metrics. Q2E is also generalizable across different text-video encoders. To show this, we applied it to the Rank 1 encoder (Wang et al., 2025) from the Zero-Shot MSR-VTT leaderboard [^5] https://paperswithcode.com/sota/zero-shot-video-retrieval-on-msr-vtt as well as the MultiCLIP encoder (following Sanders et al. (2023)).
Table 1 shows the generality of our approach. When applied to the InternVideo2, our method improved the MSR-VTT NDCG score by 1 point without audio and 4 points with audio. Additionally, the significant performance drop observed for InternVideo2 on the MultiVENT dataset underscores the dataset’s richness and complexity. Despite this drop, incorporating our method boosted the NDCG score by 19 points without audio and 26 points with audio, further validating its effectiveness.
Model R@10 P@10 MRR NDCG
Arabic
MC 72.31 62.55 0.94 76.10
MC + Q2E 74.46 63.92 0.94 78.09
MC + Q2E + ASR 79.41 68.24 0.94 82.07
Chinese
MC 74.22 68.65 0.91 77.29
MC + Q2E 82.13 76.35 0.99 86.32
MC + Q2E + ASR 83.02 76.92 0.97 86.61
English
MC 85.79 81.92 1.00 89.98
MC + Q2E 86.45 82.50 1.00 90.57
MC + Q2E + ASR 87.80 83.85 1.00 91.43
Korean
MC 65.33 62.12 0.89 70.38
MC + Q2E 72.65 69.04 0.93 76.92
MC + Q2E + ASR 76.90 73.27 0.92 80.58
Russian
MC 79.09 71.15 0.96 82.84
MC + Q2E 81.61 73.46 0.96 84.62
MC + Q2E + ASR 85.40 76.54 0.99 88.70
Table 2: Comparison of different languages. Results are reported on the MultiVENT dataset. MC = MultiCLIP encoder. Results are reported on the MultiVENT dataset. The full table is reported on Table 9 in § A.1.4.
Conversely, MultiCLIP achieved the best baseline score on the MultiVENT dataset (Sanders et al., 2023). Applying our method to MultiCLIP resulted in an 8-point NDCG improvement, setting a new state-of-the-art for the MultiVENT dataset. However, unlike InternVideo2, MultiCLIP does not use a cross-encoder reranker, which explains its comparatively lower performance on the MSR-VTT dataset. Our method still led to a 4-point improvement in the MSR-VTT NDCG score, demonstrating effectiveness across different encoders.
Next, we examine the Q2E’s performance across languages (§ 5.3), the effect of different LLM sizes (§ 5.4), different rank fusion methods (§ 5.5) and various components of our method (§ 5.6). In the appendix, we examine the impact of VLM sizes (§ A.1.1), number of frames (§ A.1.2), and frame selection methods (§ A.1.3); Q2E is quite stable across encoding decisions.
5.3 Performance Across Languages
MultiVENT includes five languages – Arabic, Chinese, English, Korean, and Russian. Table 2 shows that building upon the MultiCLIP baseline, incorporating our method consistently improves retrieval performance across all metrics. Furthermore, adding audio features results in additional gains across all languages, demonstrating the effectiveness of our approach regardless of language.
5.4 Effect of LLM size
LLM Size R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
1B 75.45 69.54 0.93 79.34
3B 75.50 69.54 0.94 79.78
8B 75.02 69.11 0.94 79.41
70B (Q2E) 75.76 69.73 0.95 80.04
With Audio
1B 78.71 72.28 0.95 82.50
3B 79.17 72.74 0.95 83.03
8B 79.04 72.59 0.95 82.91
70B (Q2E) 79.60 73.09 0.95 83.24
Table 3: Comparison of different LLM sizes within the Llama-3 family (1B to 70B parameters), specifically used for query decomposition and refining frame captions to video caption. Results are reported on the MultiVENT dataset. Full results in Table 10, § A.1.5.
Our primary model used the Llama-3.3-70B LLM. While this model is highly robust and generalizable, its large size demands significant computational power and increased inference time. To demonstrate the robustness of our method, Table 3 presents results for different LLM sizes, ranging from 1B to 70B parameters. Although the best performance is typically achieved with the 70B model, even the 1B model consistently outperforms the baseline NDCG by at least 4 NDCG points without audio and 8 points with audio. These results highlight the robustness of our method.
5.5 Effect of Rank Fusion Approaches
Method R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
Neg. Exp. Ent. 60.05 55.60 0.90 66.61
RRF 62.91 58.30 0.92 69.74
Max 71.98 66.41 0.93 76.60
Mean 73.54 67.88 0.94 78.37
Inv. Ent. (Q2E) 75.76 69.73 0.95 80.04
With Audio
Neg. Exp. Ent. 67.23 61.97 0.93 73.20
RRF 70.91 65.29 0.93 76.29
Max 76.10 70.04 0.93 80.04
Mean 78.64 72.47 0.95 82.44
Inv. Ent. (Q2E) 79.60 73.09 0.95 83.24
Table 4: Comparison of aggregation methods for zero-shot rank fusion: Neg. Exp. Ent. (Negative Exponential Entropy), RRF (Reciprocal Rank Fusion), Inv. Ent. (Inverse Entropy), Mean and Max. Results are reported on the MultiVENT dataset. Full results in Table 11, § A.1.6.
Fusing different ranks in a zero-shot setting is a non-trivial task. Traditional approaches often involve training dense networks to combine scores (Yu et al., 2024) or leveraging LLMs to ensemble multiple expert outputs (Lu et al., 2024b, a; Jiang et al., 2023b). However, rank or score fusion is a significant challenge for LLMs, as they must process and compare a large volume of numerical data, including floating-point values (for scores) and discrete values (for ranks). In both cases, LLMs struggle to produce reasonable fusion ranks. We considered several aggregation methods, from simple techniques like Mean to more complex approaches such as Reciprocal Rank Fusion (RRF) and Inverse Entropy; due to space, we describe these formulations in § A.1.6. Results in Table 4 indicate that even the simple Mean function performs well compared to RRF and Negative Exponential Entropy. Inverse Entropy demonstrated better performance, leading us to adopt it as the aggregation method.
5.6 Effect of Different Components
Components R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
Q2E 75.76 69.73 0.95 80.04
Q2E  – Video 58.00 53.90 0.89 64.83
Q2E  – Query 74.74 68.80 0.93 78.78
Q2E  – Events 74.77 68.88 0.93 79.02
With Audio
Q2E 79.60 73.09 0.95 83.24
Q2E  – Video 67.74 62.43 0.93 73.96
Q2E  – Query 78.12 71.74 0.93 81.54
Q2E  – Events 77.77 71.47 0.94 81.75
Table 5: Analysis of different components in our method. The first row in each block corresponds to our full method with all components included. In subsequent rows, one main component (video, query, event) is removed at a time to demonstrate its contribution to the overall method. Events = (Prequel + Current + Sequel). Results are reported on the MultiVENT dataset. The full result is reported on Table 12 in § A.1.7.
Table 5 presents an ablation study of query event decomposition and multimodal description components of Q2E. The results highlight that video is the most critical component. Removing the query results in a 1-point performance drop for both variants, while removing events reduces performance by 1 point in the non-audio variant and 2 points in the audio variant, highlighting the importance of decomposed events for retrieving relevant videos.
@@ 6 Conclusion & Future Work
We leveraged the world knowledge embedded in LLMs and VLMs to decompose both queries and videos for the video retrieval task, demonstrating how to match this knowledge against visual and audio signals. We employed inverse entropy as a rank fusion method without fine-tuning any additional fusion network. Experimental results across two datasets with different characteristics demonstrate the robustness and generalizability of our approach. While we show how the preexisting knowledge of LLMs and VLMs can be effectively transferred to text-to-video retrieval, our work sets the stage for future research to explore how factual and counterfactual information can be leveraged to positively or negatively align with the query.
@@ 7 Limitations
8 Acknowledgement
@@ Appendix A Appendix
A.1 Ablation Study (cont’d)
A.1.1 Effect of VLM Sizes

### v2

@@ 5 Conclusion & Future Work
We leveraged the world knowledge embedded in LLMs and VLMs to decompose both queries and videos for the video retrieval task, demonstrating how to match this knowledge against visual and audio signals. We employed inverse entropy as a rank fusion method without fine-tuning any additional fusion network. Experimental results across three datasets with different characteristics demonstrate the robustness and generalizability of our approach. While we show how the preexisting knowledge of LLMs and VLMs can be effectively transferred to text-to-video retrieval, our work sets the stage for future research to explore how factual and counterfactual information can be leveraged to positively or negatively align with the query.
@@ 6 Limitations
The primary focus of our method is to demonstrate how query enrichment through decomposition can improve understanding of human intent and ground specific events to video content. As we have not applied any enhancements to the video captions or audio transcripts, video-to-text retrieval is outside the scope of this work.
7 Acknowledgment
@@ Appendix A Appendix
A.1 Comparison with Popular Baselines
To show the robustness of our zero-shot method, in the Table˜6, we have reported the comparison of our method with different recent models: Frozen in Time (Bain et al., 2022), ImageBind (Girdhar et al., 2023), CLIP (Radford et al., 2021), CLIP4CLIP (Luo et al., 2021), CLIP2Video (Fang et al., 2021), X-Pool (Gorti et al., 2022), TS2-Net (Liu et al., 2022), TDB+TAB (Fang et al., 2023), IVRC (Tian et al., 2024), Cap4Video (Wu et al., 2023), DRL (Wang et al., 2022), T2VIndexer (Li et al., 2024), PIDRo (Guan et al., 2023b), CLIP-VIP (Xue et al., 2023), huge (Jiang et al., 2022).
[…]
MSR-VTT
Frozen in Time ICCV ’21 $\times$ 27.30 - 68.10
ImageBind CVPR ’23 $\times$ 36.80 61.80 70.00
CLIP PMLR ’21 $\times$ 39.70 72.30 82.20
CLIP4CLIP arXiv ’21 $\checkmark$ 44.50 71.40 81.60
CLIP2Video arXiv ’21 $\checkmark$ 45.60 72.60 81.70
X-Pool CVPR ’22 $\checkmark$ 46.90 72.80 82.20
TS2-Net ECCV ’22 $\checkmark$ 45.60 74.40 82.70
TDB+TAB TMM ’23 $\checkmark$ 45.60 72.60 81.70
IVRC Patt. Recog. ’24 $\checkmark$ 47.00 75.00 82.40
Cap4Video CVPR ’23 $\checkmark$ 51.40 75.70 83.90
DRL arXiv ’22 $\checkmark$ 53.30 80.30 87.60
T2VIndexer ACM MM ’24 $\checkmark$ 55.10 77.20 85.00
PIDRo ICCV ’23 $\checkmark$ 55.90 79.80 87.60
Q2E - $\times$ 56.28 76.58 83.72
CLIP-ViP ICLR ’23 $\checkmark$ 57.70 80.50 88.20
huge arXiv ’22 $\checkmark$ 62.90 84.50 90.80
MSVD
Frozen in Time ICCV ’21 $\times$ 33.70 64.70 76.30
CLIP PMLR ’21 $\times$ 37.00 64.10 73.80
CLIP4CLIP arXiv ’21 $\checkmark$ 46.20 76.10 84.60
TDB+TAB TMM ’23 $\checkmark$ 46.90 76.90 86.10
CLIP2Video arXiv ’21 $\checkmark$ 47.00 76.80 85.90
X-Pool CVPR ’22 $\checkmark$ 47.20 77.40 86.00
PIDRo ICCV ’23 $\checkmark$ 47.50 77.50 86.00
DRL arXiv ’22 $\checkmark$ 48.30 79.10 87.30
Cap4Video CVPR ’23 $\checkmark$ 51.80 80.80 88.30
T2VIndexer ACM MM ’24 $\checkmark$ 55.40 81.90 88.50
huge arXiv ’22 $\checkmark$ 59.00 84.00 90.30
Q2E - $\times$ 63.77 85.30 89.99
Table 6: Results from different baseline models reported in the recent papers. Rows are sorted based on R@1 score. FT represents the fine-tuned models for text-to-video retrieval task.
On the MSR-VTT dataset, our method outperforms all of the zero-shot methods by high margin. However, while fine-tuning can achieve moderately higher performance, our zero-shot method performs competitively. On the other hand, in the MSVD dataset, our method outperforms all baseline models, whether fine-tuned or not. This again underlines the generalizability and robustness of our method across different datasets compared to the previous baselines.
A.2 Qualitative Analysis
We have reported two examples from MSRVTT (Xu et al., 2016) and MultiVENT (Sanders et al., 2023) dataset on the Table˜7. We observed that while event decomposition may not be as grounded or clearly delineated as in the MultiVENT dataset (due to the inherent ungrounded nature of MSRVTT queries), our approach still enriches the query by introducing events that could occur before, during or after the main event. This enrichment significantly enhances the retrieval process.
Dataset MSR-VTT
Query Crocodiles about to eat the group of people walking across the water
Prequel People wading through the water, unaware of the crocodiles before they attack Crocodiles swimming stealthily towards the group during an attack on people The group noticing the crocodiles and trying to hurry across the water Crocodiles rising out of the water, jaws open People walking cautiously across the water to avoid crocodiles attacking them
Current Crocodiles attacking people while swimming towards a group of them Crocodiles emerging from the water with their mouths open during attacks on people People screaming and trying to run or swim away from crocodiles during attacks on humans People trying to hurry or run across the water when crocodiles attack them Crocodiles snapping their jaws near people during attacks
Sequel Crocodiles attacking people while chasing after a group People defending themselves against crocodile attacks when crocodiles attack them Crocodiles dragging the people underwater
Dataset MultiVENT
Query November 30 earthquake in South Central Alaska 2018
Prequel Buildings shaking and swaying due to the earthquake in Anchorage, Alaska, on November 30, 2018 People running out of buildings and evacuating the area in Anchorage, Alaska during the earthquake on November 30, 2018 Cars stopped on the road as the earthquake strikes in Anchorage, Alaska on November 30, 2018 Debris and objects falling from shelves and ceilings during the 30 November 2018 Anchorage, Alaska earthquake Emergency responders rushing to the scene to assist with evacuation and relief efforts after the 30 November 2018 earthquake in Anchorage, Alaska, South Central Alaska
Current Buildings shaking and crumbling during the 30 November 2018 earthquake in Anchorage, Alaska, South Central Alaska People running out of buildings and evacuating the area during the 30 November 2018 Anchorage, Alaska earthquake Emergency responders rushing to the scene after the 2018 Anchorage, Alaska earthquake on November 30, 2018 Debris falling from buildings and damaging streets during the November 30, 2018 earthquake in Anchorage, South Central Alaska Cars stopped or abandoned on the road in Anchorage, Alaska, South Central Alaska due to the earthquake on November 30, 2018
Sequel Buildings crumbling or collapsing during the November 30, 2018 earthquake in Anchorage, Alaska People running for cover or evacuating buildings during the November 30, 2018 earthquake in Anchorage, Alaska Emergency vehicles rushing to the scene of the 2018 Anchorage, Alaska earthquake on November 30, 2018 Cracks forming in roads and highways in Anchorage, Alaska, South Central Alaska, after the 30 November 2018 earthquake Debris falling from damaged structures during the 30 November 2018 earthquake in Anchorage, Alaska
Table 7: Qualitative analysis on two examples from MSRVTT (Xu et al., 2016) and MultiVENT(Sanders et al., 2023) dataset respectively.
From the retrieval ranking perspective, incorporating event-based decomposition substantially improves video retrieval accuracy. For example, for the provided MSR-VTT query, the ground-truth video is ranked 12th (with audio) and 13th (without audio) by MultiCLIP, whereas Q2E (which integrates decomposition module, multimodal features, textual descriptions, and our ASR module) elevates it to 1st place with audio and 7th place without audio.
A.3 Ablation Study (continued)
A.3.1 Effect of VLM Sizes